OpenStudy (anonymous):

Identify the number of solutions for the systems of equations 3x-12y=6 4y=12-6x a) one b)none c)infinite d) cannot be determined

OpenStudy (amistre64):

lets get them looking the same; get all variables to one side and constants to the other like this: 3x -12y = 6 6x +4y = 12

OpenStudy (amistre64):

if they are parallel, they have the same slope: -3/-12 = -6/4 not gonna be parallel so (c) is out of the picture

OpenStudy (amistre64):

since they are lines; they can be determined to (d) is out of the picture as well

OpenStudy (amistre64):

they are not the same line, or else they d be parallel ; and they aint parallel so (c); as well as (b) is out

OpenStudy (amistre64):

go with (a)

OpenStudy (anonymous):

thanks